One third of a number is -92 Find the number
Answer by ptaylor(2198)

Let x=the number
(1/3)x=-92 multiply each side by 3
3*(1/3)x=-92*3
x=-276------------------the number
CK
1/3 of -276=-92

Hope this helps---ptaylor